# Inferring Thunderstorm Occurrence from Vertical Profiles of Convection-Permitting Simulations: Physical Insights from a Physical Deep Learning Model

### v1

@@ 5 Discussion and conclusion
Bypassing the traditional use of derived single-level predictors from NWP data, we developed SALAMA 1D, an ML model for predicting the probability of thunderstorm occurrence on a pixel-wise basis by processing vertical profiles of three-dimensional variables from convection-permitting NWP forecasts. The model’s architecture was motivated by physical considerations. In particular, a sparse layer reduced parameter size by encouraging interactions at similar height levels, while a shuffling mechanism prevented the model from learning patterns tied to the vertical grid structure. The latter also added a form of regularization that limited overfitting.
@@
In comparison to SALAMA 0D, an ML baseline model that infers thunderstorm occurrence from derived single-level features, our model demonstrated higher skill across a wide range of metrics and for lead times up to at least . This result indicates that information relevant to thunderstorm occurrence, while intricately encoded in vertical profiles, can be successfully extracted by machine learning, resulting in an improved ability to recognize thunderstorm occurrence in NWP forecasts. Notably, our model achieves this enhanced skill with only a modest increase in computational time during prediction, making it equally suitable for real-time operational use as SALAMA 0D. Furthermore, doubling the number of days used to compile the training set (while keeping the training set size constant) also increased skill, underscoring the importance of a large and diverse database of NWP data. We anticipate further skill improvements with the collection of more NWP data.
@@
A sensitivity analysis based on saliency maps indicated that our model prioritizes temperature and pressure to reconstruct environmental lapse rates, possibly in order to infer levels of positive buoyancy and convection inhibition. Additional information is gathered from ice particle content near the tropopause, while the absence of mid-level graupel and cloud cover is used to rule out thunderstorm occurrence. The model relies significantly on mesoscale structures related to convective initiation, such as lapse rates and low-level moisture, and to a lesser extent on fine-grained structures characteristic of ongoing convection. We hypothesized that this preference arises because larger-scale patterns are more reliable for identifying thunderstorm occurrence, as fine-grained patterns may not consistently align with lightning observations due to the convection-permitting NWP model struggling to accurately predict the location and timing of convection. To encourage the use of fine-grained structures, it may be beneficial to train the model in such a way that it can process all ensemble members simultaneously in one forward pass. By analyzing the number of members displaying fine-grained signs of thunderstorm activity, the model could learn the degree to which it can trust the convection produced by NWP.
@@ Acknowledgements.
- Markowski and Richardson (2011) Markowski, P., and Y. Richardson, 2011: Mesoscale meteorology in midlatitudes. John Wiley & Sons, 42-43 pp.
@@
- Ukkonen and Mäkelä (2019) Ukkonen, P., and A. Mäkelä, 2019: Evaluation of machine learning classifiers for predicting deep convection. Journal of Advances in Modeling Earth Systems, 11 (6), 1784–1802, https://doi.org/10.1029/2018MS001561, %**** ̵main.bbl ̵Line ̵275 ̵****https://agupubs.onlinelibrary.wiley.com/doi/pdf/10.1029/2018MS001561.

### v2

@@ 5 Discussion and conclusion
Bypassing the traditional use of derived single-level predictors from NWP data, we developed SALAMA 1D, an ML model for predicting the probability of thunderstorm occurrence on a pixel-wise basis by processing vertical profiles of three-dimensional variables from convection-permitting NWP forecasts. The model’s architecture was motivated by physical considerations. In particular, a sparse layer reduces parameter size by encouraging interactions at similar height levels, while a shuffling mechanism prevents the model from learning patterns tied to the vertical grid structure. The latter also adds a form of regularization that limits overfitting.
@@
In comparison to SALAMA 0D, an ML baseline model that infers thunderstorm occurrence from derived single-level features, our model demonstrated higher skill across a wide range of metrics and for lead times up to at least . This result indicated that information relevant to thunderstorm occurrence, while intricately encoded in vertical profiles, can be successfully extracted by machine learning, resulting in an improved ability to recognize thunderstorm occurrence in NWP forecasts. Notably, our model remains lightweight in terms of computational complexity, making it just as suitable for real-time operational use as SALAMA 0D is. Case studies suggested that SALAMA 1D is capable of correcting the raw NWP output when convective areas are of incorrect size or when NWP fails to produce convection in the first place. Furthermore, doubling the number of days used to compile the training set (while keeping the training set size constant) also increased skill, underscoring the importance of a large and diverse database of NWP data. We anticipate further skill improvements with the collection of more NWP data.
@@
A sensitivity analysis based on saliency maps revealed that many learned patterns are physically interpretable and indicated that our model significantly relies on mesoscale patterns: For instance, our results suggested that SALAMA 1D has learned the climatological propagation direction of thunderstorms in the study region and involves temperature, pressure and moisture profiles to infer atmospheric instability and, hence, the potential for convective initiation. To a lesser extent, the model relies on more fine-grained structures characteristic of ongoing convection, such as ice particle content near the tropopause and cloud cover. We hypothesized that the preference for mesoscale patterns arises because of their higher reliability for identifying thunderstorm occurrence, as fine-grained patterns may not consistently align with lightning observations due to the convection-permitting NWP model struggling to accurately predict the location and timing of convection. To encourage the use of fine-grained structures, it may be beneficial to train the model in such a way that it can process all ensemble members simultaneously in one forward pass. By analyzing the number of members displaying fine-grained signs of thunderstorm activity, the model could learn the degree to which it can trust the convection produced by NWP. Furthermore, one could adapt the model in such a way that horizontally extended input, or several forecast times around the target time, are processed. This would allow for improvement on correcting location and timing errors of convection.
@@ Acknowledgements.
- Carlson et al. (1983) Carlson, T. N., S. G. Benjamin, G. S. Forbes, and Y.-F. Li, 1983: Elevated mixed layers in the regional severe storm environment: Conceptual model and case studies. Monthly Weather Review, 111 (7), 1453 – 1474, 10.1175/1520-0493(1983)111¡1453:EMLITR¿2.0.CO;2.
@@
- Dramsch et al. (2025) Dramsch, J. S., and Coauthors, 2025: Explainability can foster trust in artificial intelligence in geoscience. Nature Geoscience, 1–3.
@@
- Hubbert et al. (2018) Hubbert, J. C., J. W. Wilson, T. M. Weckwerth, S. M. Ellis, M. Dixon, and E. Loew, 2018: S-pol’s polarimetric data reveal detailed storm features (and insect behavior). Bulletin of the American Meteorological Society, 99 (10), 2045 – 2060, 10.1175/BAMS-D-17-0317.1.
@@
- Leinonen et al. (2022) Leinonen, J., U. Hamann, U. Germann, and J. R. Mecikalski, 2022: Nowcasting thunderstorm hazards using machine learning: the impact of data sources on performance. Natural Hazards and Earth System Sciences, 22 (2), 577–597, 10.5194/nhess-22-577-2022.
@@
- Markowski and Richardson (2010) Markowski, P., and Y. Richardson, 2010: Basic Equations and Tools. John Wiley & Sons, Ltd, https://doi.org/10.1002/9780470682104, URL https://onlinelibrary.wiley.com/doi/abs/10.1002/9780470682104, https://onlinelibrary.wiley.com/doi/pdf/10.1002/9780470682104.
- Marquis et al. (2023) Marquis, J. N., Z. Feng, A. Varble, T. C. Nelson, A. Houston, J. M. Peters, J. P. Mulholland, and J. Hardin, 2023: Near-cloud atmospheric ingredients for deep convection initiation. Monthly Weather Review, 151 (5), 1247 – 1267, 10.1175/MWR-D-22-0243.1.
@@
- Mueller et al. (2003) Mueller, C., T. Saxen, R. Roberts, J. Wilson, T. Betancourt, S. Dettling, N. Oien, and J. Yee, 2003: Ncar auto-nowcast system. Weather and Forecasting, 18 (4), 545 – 561, https://doi.org/10.1175/1520-0434(2003)018¡0545:NAS¿2.0.CO;2.
@@
- Ortland et al. (2023) Ortland, S. M., M. J. Pavolonis, and J. L. Cintineo, 2023: The development and initial capabilities of thundercast, a deep learning model for thunderstorm nowcasting in the united states. Artificial Intelligence for the Earth Systems, 2 (4), e230 044, 10.1175/AIES-D-23-0044.1.
@@
- Peters et al. (2023) Peters, J. M., D. R. Chavas, C.-Y. Su, H. Morrison, and B. E. Coffer, 2023: An analytic formula for entraining cape in midlatitude storm environments. Journal of the Atmospheric Sciences, 80 (9), 2165 – 2186, 10.1175/JAS-D-23-0003.1.
@@
- Sovrasov (2018) Sovrasov, V., 2018: ptflops: a flops counting tool for neural networks in pytorch framework. URL https://github.com/sovrasov/flops-counter.pytorch.
@@
- Theis et al. (2005) Theis, S. E., A. Hense, and U. Damrath, 2005: Probabilistic precipitation forecasts from a deterministic model: a pragmatic approach. Meteorological Applications, 12 (3), 257–268, https://doi.org/10.1017/S1350482705001763, https://rmets.onlinelibrary.wiley.com/doi/pdf/10.1017/S1350482705001763.
- Tuckman et al. (2023) Tuckman, P., V. Agard, and K. Emanuel, 2023: Evolution of convective energy and inhibition before instances of large cape. Monthly Weather Review, 151 (1), 321 – 338, 10.1175/MWR-D-21-0302.1.
- Ukkonen and Mäkelä (2019) Ukkonen, P., and A. Mäkelä, 2019: Evaluation of machine learning classifiers for predicting deep convection. Journal of Advances in Modeling Earth Systems, 11 (6), 1784–1802, https://doi.org/10.1029/2018MS001561, https://agupubs.onlinelibrary.wiley.com/doi/pdf/10.1029/2018MS001561.
@@
- Vivekanandan et al. (1999) Vivekanandan, J., D. S. Zrnic, S. M. Ellis, R. Oye, A. V. Ryzhkov, and J. Straka, 1999: Cloud microphysics retrieval using s-band dual-polarization radar measurements. Bulletin of the American Meteorological Society, 80 (3), 381 – 388, 10.1175/1520-0477(1999)080¡0381:CMRUSB¿2.0.CO;2.
@@
- Zhang et al. (2003) Zhang, F., C. Snyder, and R. Rotunno, 2003: Effects of moist convection on mesoscale predictability. Journal of the Atmospheric Sciences, 60 (9), 1173 – 1185, 10.1175/1520-0469(2003)060¡1173:EOMCOM¿2.0.CO;2.